OpenStudy (anonymous):

what is a three step equation in which the answer is 24?

OpenStudy (anonymous):

\[2x+1=x+25\]

OpenStudy (anonymous):

2x + x + 1 + 1 = x + 25 + 25

OpenStudy (anonymous):

all wrong

OpenStudy (anonymous):

there is more than one correct answer ..

OpenStudy (anonymous):

give me one that the answer is 24

OpenStudy (anonymous):

2x + x + 1 + 1 = x + 25 + 25

OpenStudy (anonymous):

no cuz i solved it and i got 12 so Luke can u please solve it so i can see what i did wrong.

OpenStudy (anonymous):

:-/

OpenStudy (anonymous):

2x + x + 1 + 1 = x + 50 + 50

OpenStudy (anonymous):

Love the gratitude for freely volunteered time I must say.

OpenStudy (anonymous):

all wrong. lol

OpenStudy (anonymous):

.... give me the answer ...

OpenStudy (anonymous):

yea the answer is 49 and i am looking for an equation that the answer is 24 not 49

OpenStudy (anonymous):

apparently the method "make up an expression, replace x by 24 and see what you get" does not work. do you have something specific in mind?

OpenStudy (anonymous):

1+2x=49?

OpenStudy (anonymous):

1+x=-x+49?

OpenStudy (anonymous):

can someone give me another one please?

OpenStudy (anonymous):

\[\Large \begin{array}{l} 2x{\rm{ }} + {\rm{ }}x{\rm{ }} + {\rm{ }}1{\rm{ }} + {\rm{ }}1{\rm{ }} = {\rm{ }}x{\rm{ }} + {\rm{ }}25{\rm{ }} + {\rm{ }}25\\ 2x = 48\\ x = \frac{{48}}{2}\\ x = 24 \end{array}\] Jaysus. Good answer Luke.

OpenStudy (anonymous):

how u get 48

OpenStudy (anonymous):

Collect like terms and simplify. The 1 + 1 is one one side, move it to the other. 25 + 25 = 50 50 - (1+1) = 48. \[\Large \begin{array}{l} 2x{\rm{ }} + {\rm{ }}x{\rm{ }} + {\rm{ }}1{\rm{ }} + {\rm{ }}1{\rm{ }} = {\rm{ }}x{\rm{ }} + {\rm{ }}25{\rm{ }} + {\rm{ }}25\\ 2x + x - x = 25 + 25 - 1 - 1 \end{array}\]

OpenStudy (anonymous):

And just for the record, being a little less demanding and a little more appreciative of people's freely-offered help would be a decent gesture.

OpenStudy (anonymous):

can u just give me another one becuz i don't understand that one

OpenStudy (anonymous):

\[\Large \begin{array}{l} 2x + 4 = 52\\ 2x = 52 - 4 = 48\\ x = \frac{{48}}{2}\\ x = 24 \end{array}\]

OpenStudy (anonymous):

ok thank u and how about an equation where the answer is 26